# MMS-LLaMA: Efficient LLM-based Audio-Visual Speech Recognition with Minimal Multimodal Speech Tokens

### v1

@@ 6 Conclusion
We have demonstrated that the proposed MMS-LLaMA, integrating efficient token compression strategies into multimodal speech LLM frameworks can dramatically reduce computational costs while preserving high-level linguistic content. By incorporating an early AV-fusion module, a dynamically adaptive AV Q-Former, and a refined query allocation strategy with a speech rate predictor, we have achieved state-of-the-art AVSR performance, attaining a WER of 0.74% while using only 3.5 multimodal speech tokens per second. Moreover, our extensive experiments on the LRS3 dataset have confirmed that this method not only achieves a superior WER but also reduces FLOPs by 35.7% compared to the previous LLM-based AVSR approach.

### v2

@@ 6 Conclusion
We have demonstrated that the proposed MMS-LLaMA, integrating efficient token compression strategies into multimodal speech LLM frameworks can dramatically reduce computational costs while preserving high-level linguistic content. By incorporating an early AV-fusion module, a dynamically adaptive AV Q-Former, and a refined query allocation strategy with a speech rate predictor, we have achieved state-of-the-art AVSR performance, attaining a WER of 0.72% while using only 3.5 multimodal speech tokens per second. Moreover, our extensive experiments on the LRS3 dataset have confirmed that this method not only achieves a superior WER but also reduces FLOPs by 35.7% compared to the previous LLM-based AVSR approach.
Method WER (%)
VSP-LLM Yeo et al. (2024a) 26.7
Proposed Method 28.7
Table 7: Performance comparison for the visual-only modality on LRS3.
Method SNR (dB)
-15 -10 -5 0 5 10 15
Cheng et al. (2023a) 34.9 16.6 5.8 2.6 2.0
Proposed Method 17.1 13.8 7.0 1.9 1.0 0.76 0.75
Table 8: WER (%) of different methods at various SNR levels (dB).
@@ 7 Limitation
Acknowledgments
This work was partly supported by two funds: the National Research Foundation of Korea (NRF) grant funded by the Korea government (MSIT) (No. NRF-2022R1A2C2005529), and Institute of Information & communications Technology Planning & Evaluation (IITP) grant funded by the Korea government(MSIT) (No.2022-0-00124, Development of Artificial Intelligence Technology for Self-Improving Competency-Aware Learning Capabilities).